OpenStudy (anonymous):

solve equation in the interval from 0 to 2pi. 7 cos 2t =3

OpenStudy (anonymous):

corrections 7 cos 2t=3

OpenStudy (cwrw238):

arccos 3/7 - your calculator will give u the value of t in first quadrant the other value is in 4th quadrant

OpenStudy (anonymous):

what does arccos mean

OpenStudy (anonymous):

am I setting calculator in rad. or degrees

OpenStudy (cwrw238):

oh - then find angle whose cos is 3/7 0 calculator will give you value of 2t in first quadrant

OpenStudy (cwrw238):

right - set to rad arccos means the angle whose cos is given value or - on calculator button this might be cos -1

OpenStudy (anonymous):

cos 3/7= .91 now what

OpenStudy (anonymous):

how do I get the other value

OpenStudy (cwrw238):

you need cos-1 not cos check again : i make it 1.128 rad

OpenStudy (anonymous):

ok that is 1.099??

OpenStudy (anonymous):

1/cos(3/7) right?

OpenStudy (anonymous):

and how do you know I need cos-1 and not cos?

OpenStudy (cwrw238):

no - the cosine of the angle 2t = 3/7 you need to find the angle whose cosine is 3/7 first find 3 / 7 on ur calculator then you need to press the cos-1 (or arccos) button

OpenStudy (anonymous):

do I just press cos then -1

OpenStudy (cwrw238):

cos-1 does not mean 1 / cos in this context

OpenStudy (cwrw238):

no there is probably a button marked cos but you first press the shift key then cos button to get value of the angle

OpenStudy (cwrw238):

cos-1 notation is a bit confusing - in Europe they use arccos which makes more sense

OpenStudy (anonymous):

cos (3/7)=.91. Do I press shift key before entering cos (3/7)? don't know the shift key

OpenStudy (anonymous):

I haven't done any problems with cos-1

OpenStudy (cwrw238):

calculators vary you either press shift cos (3/7) or (3/7) shift cos

OpenStudy (cwrw238):

the correct value is 1.127 (to 3 dec places)

OpenStudy (cwrw238):

sorry 1.128

OpenStudy (cwrw238):

its a bit hard to explain without seeing your calculator....